OpenStudy (anonymous):

A repulsive force of 14.4 N occurs between two charges that are 0.10 m apart. If one charge is -4.0 μC, what is the other charge? -1.0 μC -4.0 μC +1.0 μC +2.0 μC

OpenStudy (irishboy123):

use Coulumb's Law: \(F = \frac{k\ q_1\ q_1 }{r^2}\), where: \( k = 9×10^9 Nm^2/C^2 \) force is repulsive so other charge must be .......

OpenStudy (anonymous):

-4.0 μC ?